# Solving rescheduling problems in heterogeneous urban railway networks using hybrid quantum-classical approach

### v4

@@ 4 Methods and Model
4.1 Sets, parameters and decisions
@@
To formulate our decision variables, constraints, and the objective function, we determine sets of index tuples needed to find the actual index sets of variables, from the given infrastructure data, timetable data, and the rolling stock circulation plan. Also, we introduce parameters calculated from the same input.
Sets
Let us denote by $\hat{S}$ the set of all stations, and by $S\subset\hat{S}$ the decision stations we have determined in advance, as already described. The main objects of our model for railway rescheduling are the trains $j\in\mathcal{J}$ and the stations $s\in\mathcal{S}_{j}$ in their path. Set $\mathcal{S}_{j}$ includes decision stations only and it is an ordered set. In addition to $\mathcal{J}$ and $\mathcal{S}_{j}$, the relevant sets are the following. Set $\mathcal{J}^{2\,(\text{turn})}_{s}\subset\mathcal{J}\times\mathcal{J}\left(\forall s\in\mathcal{S}\right)$ is the set of all pairs of trains so that the first train of the pair terminates at station $s$ and its rolling stock continues as the second train of the pair. This set is deduced from the rolling stock circulation plan and the timetable. Set $\mathcal{J}^{2\,(\text{close})}\subset\mathcal{J}\times\mathcal{J}$ is the set of trains that are close enough to each other in time so that precedence variables have to be defined for them. More details will be given in the description of the parameter $d_{\max}$ which this set depends on. Set $\mathcal{J}^{2\,(\text{headway})}\subset\mathcal{J}^{2\,(\text{close})}$ is the set of train pairs that can potentially violate the minimal headway time between two subsequent trains on a line segment. For a pair of trains to be in $\mathcal{J}^{2\,(\text{headway})}$, both their routes need to include the same line segment so that the trains are moving in the same direction on it and can meet there according to model parameters, i.e. maximal allowed delay. Set $\mathcal{J}^{2\,(\text{single})}\subset\mathcal{J}^{2\,(\text{close})}$ is the set of all trains that share at least one single-track line segment as a common part of their routes so that they are heading in the opposite direction. Set $\left(\forall s\in\mathcal{S}\right)\ \mathcal{J}^{2\,(\text{track})}_{s}\subset\mathcal{J}^{2\,(\text{close})}$ is the set of train pairs that are planned to occupy the same track on station $s$ anytime during the planning horizon, thereby competing for the same track. Set $\left(\forall s\in\mathcal{S}\right)\ \mathcal{J}^{2\,(\text{switch},\text{out})}_{s}\subset\mathcal{J}^{2\,(\text{close})}$ is the set of train pairs that are planned to pass the same interlocking area of $s$ upon their departure from station $s$. Set $\left(\forall(s,s^{\prime})\in\mathcal{S}^{\times 2}\right)\ \mathcal{J}^{2\,(\text{switch},\text{out},\text{in})}_{s,s^{\prime}}\subset\mathcal{J}^{2\,(\text{close})}$ is the set of train pairs that are planned to pass the same interlocking area of $s$ in order to have $j$ to departure $s$ while $j^{\prime}$ arrive $s$ from the direction of $s^{\prime}$. Set $\left(\forall s\in\mathcal{S}^{\times 2}\right)\ \mathcal{J}^{2\,(\text{switch},\text{in},\text{noMP})}_{s,s^{\prime}}\subset\mathcal{J}^{2\,(\text{close})}$ is the set of train pairs that are planned to pass the same interlocking area of $s$ upon their arrival at $s$ from the direction of $s^{\prime}$, and there is no M-P possibility for them between $s$ and $s^{\prime}$ Set $\left(\forall s\in\mathcal{S}^{\times 2}\right)\ \mathcal{J}^{2\,(\text{switch},\text{in},\text{MP})}_{s,s^{\prime}}\subset\mathcal{J}^{2\,(\text{close})}$ is the set of train pairs that are planned to pass the same interlocking area of $s$ upon their arrival at $s$ so that either they both come from the direction of $s^{\prime}$ but there is a M-P possibility for them between $s$ and $s^{\prime}$, or one of them is approaching $s$ from a direction other than $s^{\prime}$. Set $\left(\forall j\in\mathcal{J}\right)\ \mathcal{C}^{2}_{j}$ is the set of all station pairs that are subsequent in the route of $j$. Set $\left(\forall(j,j^{\prime})\in\mathcal{J}^{\times 2}\right)\ \mathcal{C}^{2\,(\text{common})}_{j,j^{\prime}}$ is the set of all station pairs that appear as subsequent stations in the route of both $j$ and $j^{\prime}$ heading in the same direction. Set $\left(\forall(j,j^{\prime})\in\mathcal{J}^{\times 2}\right)\ \mathcal{C}^{2\,(\text{common, single})}_{j,j^{\prime}}$ is the set of all station pairs that appear as subsequent stations in the route of both $j$ and $j^{\prime}$ which are connected with a single-track line segment, and trains are heading in opposite directions. The order within the pairs is determined by $j$. All these sets can be enumerated based on the input data in a straightforward manner.
Parameters
@@
The parameters that appear in our model are the following:
@@
  • $\tau^{(\text{switch})}(j,j^{\prime},s)$ is the running time of $j$ over a interlocking area of $s$, where $j$ may be in conflict with $j^{\prime}$. (In our examples we do not consider cases when, e.g., on bigger stations there are multiple switches with different technological times. The model could trivially be extended to cover such scenarios by adding extra indices if relevant.)
@@
  • $\sigma(j,s)$ is the scheduled departure time of $s$ from $j$, given by the original timetable.
  • $\upsilon(j,s)$ is the earliest possible departure time of $j$ from $s$ when a disturbance in the network happens. It is the maximum of $\sigma(j,s)$ and the technically feasible earliest departure time, and may be more constraining than the first depending on the initial delays. The latter is calculated assuming that the train follows its planned route at minimum running time and that there are no other trains present on the network.
@@
  • $d_{\text{max}}$ is an upper bound assumed for the secondary delay $t^{(\text{out})}(j,s)-\upsilon(j,s)$. This parameter sets an upper limit for the possible secondary delays that can arise on a particular part of the network. Setting such a bound is common in the literature, see e.g. the work of D’Ariano et al. (2007). We set the same value of this parameter for all our testing instances. It has to be big enough so that no secondary delay exceeds it (e.g. $d_{\max}=40$ excludes the possibility of a one-hour delay due to waiting for another delayed train). Meanwhile, it is desirable to set its value as small as possible; restricting the time variables to small intervals decreases the number of binary decision variables in the model.
@@
The choice of a small $d_{\max}$ parameter results in a decrease of the model size, which is achieved via defining the set $J^{2\,(\text{close})}$. A pair of trains $(j,j^{\prime})$ is included in this set if and only if they can meet on any station, given the original timetable, the disturbed timetable, and assuming that no train can have a secondary delay greater than $d_{\max}$. The number of constraints will be proportional to the average number of trains that can meet another train at a station; this will be linear in $d_{\max}$.
Decision variables
@@
Our decision variables are the following. First, we use departure time variables:
$t^{(\text{out})}(j,s)\in\mathbb{N}$ (1)
@@
defining the departure time of train $j\in\mathcal{J}$ from station $s\in\mathcal{S}_{j}$. Such a variable is defined for all decision stations $\mathcal{S}_{j}$. In the case of trains that terminate within the modeled part of the network, the last station is not included in $\mathcal{S}_{j}$ The arrival time of the trains at stations, $t^{(\text{in})}(j,s)\in\mathbb{N}$ are trivially related to the $t^{(\text{out})}(j,s)$ variables through a constant offset (c.f. Eq. (3)), given the fixed running time assumption.
@@
Second, in addition to the time variables, we use three sets of binary precedence variables. Decision variables $y^{(\text{out})}(j,j^{\prime},s)\in\{0,1\}$ determine the order of trains to leave stations: the variable takes the value $1$ if train $j$ leaves station $s$ before train $j^{\prime}$, and 0 otherwise. The $(j,j^{\prime},s)$ tuples for which an $y$ variable is defined will be specified later. Similarly, the precedence variables $y^{(\text{in})}(j,j^{\prime},s)\in\{0,1\}$ prescribe the order at the entry to stations; the value is $1$ if $j$ arrives to $s$ before $j^{\prime}$. The last set of precedence variables describes the precedence of trains at some resource located between stations $s$ and $s^{\prime}$ (e.g. a single-track segment used by $j$ and $j^{\prime}$ heading in opposite direction). The binary variable $z(j,j^{\prime},s,s^{\prime})\in\{0,1\}$ will be $1$ if $j$ uses the given resource before $j^{\prime}$. Also in this case, the quadruples $(j,j^{\prime},s,s^{\prime})$ for which we have such a variable will be specified later.
@@ 4.2 ILP formulation
Objective function
Our goal is to minimize the secondary delays that occur in the analyzed part of the railway network. Hence, a suitable objective function is the weighted sum of secondary delays at the destination station:
$f(t)=\frac{1}{d_{\text{max}}}\sum_{j\in\mathcal{J}}w(j)\left(t^{(\text{out})}(j,s^{*})-\upsilon(j,s^{*})\right).$ (2)
where $s^{*}$ is the last element of $S_{j}$, and $w_{j}$ are weights for each train representing its priority. The reason for restricting ourselves to the secondary delays at the destination stations only is that this is the actual figure of merit chosen by the respective railway authorities; the generalization to other linear objectives such as the weighted sum of secondary delays on all stations is straightforward. The constant multiplier $1/d_{\text{max}}$ is optional; we use it for better comparability of different instances.
Constraints
@@
$\forall j\in\mathcal{J}\ \forall(s,s^{\prime})\in\mathcal{C}_{j}\quad t^{(\text{in})}(j,s^{\prime})=t^{(\text{out})}(j,s)+\tau^{(\text{pass})}(j,s\rightarrow s^{\prime})$ (3)
@@
$\forall(j,j^{\prime})\in\mathcal{J}^{2\,(\text{headway})}\ \forall(s,s^{\prime})\in\mathcal{C}^{2\,(\text{common})}_{j,j^{\prime}}t^{(\text{out})}(j^{\prime},s)\geq$
$t^{(\text{out})}(j,s)+\tau^{(\text{headway})}(j,j^{\prime},s\rightarrow s^{\prime})-C\cdot y^{(\text{out})}(j^{\prime},j,s),$ (4)
where $C$ is a constant big enough to make the constraint satisfied whenever the binary variable $y^{(\text{out})}(j^{\prime},j,s)$ takes the value of $1$. (Constraints of this structure are often termed as big-M constraints in this context.) In our implementation, we calculate and use the smallest suitable value of $C$ given particular $d_{\max}$ value, i.e.
$C=-\min(t^{(\text{out})}(j^{\prime},s))+\max(t^{(\text{out})}(j,s))+\tau^{(\text{headway})}(j,j^{\prime},s\rightarrow s^{\prime})$ (5)
@@
$\forall(j,j^{\prime})\in\mathcal{J}^{2\,(\text{single})}\ \forall(s,s^{\prime})\in\mathcal{C}^{2\,(\text{common, single})}_{j,j^{\prime}}$
$t^{(\text{out})}(j^{\prime},s^{\prime})\geq t^{\text{(in)}}(j,s^{\prime})-C\cdot z(j^{\prime},j,s^{\prime},s),$ (6)
where $C$ is a big enough constant chosen similarly to that in Eq. (4).
@@
$\forall j\in\mathcal{J}\forall s\in S_{j}\quad t^{(\text{out})}(j,s)\geq t^{(\text{in})}(j,s)+\tau^{(\text{dwell})}(j,s).\$ (7)
Timetable
No train is allowed to depart before its scheduled departure time:
$\forall j\in\mathcal{J}\forall s\in S_{j}\quad t^{(\text{out})}(j,s)\geq\sigma(j,s).$ (8)
@@
${\forall s\in S\forall(j,j^{\prime})\in\mathcal{J}^{2\,(\text{track})}_{s}}t^{(\text{in})}(j^{\prime},s)\geq t^{(\text{out})}(j,s)-C\cdot y^{(\text{out})}(j^{\prime},j,s),\$ (9)
where $C$ is chosen similarly to that in Eq. (4) again. Note that this requirement may not be needed for depot tracks; the exceptions can be handled by the proper definition of $\mathcal{J}^{2\,(\text{track})}_{s}$.
@@
$\forall s\in S\forall(j,j^{\prime})\in\mathcal{J}^{2\,(\text{switch},\text{out})}_{s}t^{(\text{out})}(j^{\prime},s)\geq$
$t^{(\text{out})}(j,s)+\tau^{(\text{switch})}(j,j^{\prime},s)-C\cdot y^{(\text{out})}(j^{\prime},j,s),$ (10)
$\forall(s,s^{\prime})\in S^{\times 2}\forall(j,j^{\prime})\in\mathcal{J}^{2\,(\text{switch},\text{out},\text{in})}_{s,s^{\prime}}t^{(\text{in})}(j^{\prime},s)\geq$
$t^{(\text{out})}(j,s)+\tau^{(\text{switch})}(j,j^{\prime},s)-C\cdot z(j^{\prime},j,s^{\prime},s),$ (11)
$\forall(s,s^{\prime})\in S^{\times 2}\forall(j,j^{\prime})\in\mathcal{J}^{2\,(\text{switch},\text{in},\text{noMP})}_{s,s^{\prime}}t^{(\text{in})}(j^{\prime},s)\geq$
$t^{(\text{in})}(j,s)+\tau^{(\text{switch})}(j,j^{\prime},s)-C\cdot y^{(\text{out})}(j^{\prime},j,s^{\prime}),$ (12)
$\forall(s,s^{\prime})\in S^{\times 2}\forall(j,j^{\prime})\in\mathcal{J}^{2\,(\text{switch},\text{in},\text{MP})}_{s,s^{\prime}}t^{(\text{in})}(j^{\prime},s)\geq$
$t^{(\text{in})}(j,s)+\tau^{(\text{switch})}(j,j^{\prime},s)-C\cdot y^{(\text{in})}(j^{\prime},j,s),$ (13)
with a choice of $C$ similar again to that in Eq. (4).

### v5

@@ 4 Methods and Model
4.1 Sets, parameters and decision variables
@@
In order to define our decision variables, constraints, and objective function, we first determine sets of index tuples from the given infrastructure data, timetable data, and the rolling stock circulation plan. Also, we introduce parameters calculated from the same input.
4.1.1 Sets
$\hat{S}$ stations
$S$ decision stations
$J$ trains
$\mathcal{S}_{j}$ stops of train $j$ (ordered)
$\mathcal{J}^{2\,(\text{close})}$ train pairs with precedence variables
$\mathcal{J}^{2\,(\text{single})}$ train pairs possibly conflicting on single-track segments
$\mathcal{J}^{2\,(\text{headway})}$ train pairs prone to violation of minimal headway
$\mathcal{J}^{2\,(\text{turn})}_{s}$ train pairs with the same rolling stock turning at $s$
$\mathcal{J}^{2\,(\text{track})}_{s}$ train pairs possibly competing for the same track of $s$
$\mathcal{J}^{2\,(\text{switch},\text{out})}_{s}$ train pairs possibly competing for an interlocking area of $s$ upon departure
$\mathcal{J}^{2\,(\text{switch},\text{out},\text{in})}_{s,s^{\prime}}$ train pairs possibly competing for an interlocking area of $s$ upon the departure of $j$ from $s$ while $j^{\prime}$ arrives to $s$ from the direction of $s^{\prime}$.
$\mathcal{J}^{2\,(\text{switch},\text{in},\text{noMO})}_{s,s^{\prime}}$ train pairs possibly competing for an interlocking area of $s$ upon arrival from $s^{\prime}$ under special circumstances described in the main text.
$\mathcal{J}^{2\,(\text{switch},\text{in},\text{MO})}_{s,s^{\prime}}$ train pairs possibly competing for an interlocking area of $s$ upon arrival, under special circumstances described in the main text.
$\mathcal{C}^{2}_{j}$ station pairs subsequent in the route of $j$
$\mathcal{C}^{2\,(\text{common})}_{j,j^{\prime}}$ station pairs subsequent in the route of both $j$ and $j^{\prime}$ heading in the same direction
$\mathcal{C}^{2\,(\text{common, single})}_{j,j^{\prime}}$ station pairs which are connected with a single-track line segment, subsequent in the route of both $j$ and $j^{\prime}$ heading in the opposite directions
Table 1: Summary of sets
Let $\hat{S}$ denote the set of all stations, and $S\subset\hat{S}$ the decision stations which have been selected in advance, as already described. The main objects of our model for railway rescheduling are the trains $j\in\mathcal{J}$ and the stations $s\in\mathcal{S}_{j}$ in their path. The sets $\mathcal{S}_{j}$ list decision stations only and they are ordered. In addition to $\mathcal{J}$ and $\mathcal{S}_{j}$, the relevant sets are the following.
  • The set $\mathcal{J}^{2\,(\text{close})}\subset\mathcal{J}\times\mathcal{J}$ is the set of train pairs for which a precedence variable can potentially be defined in the model. It could be considered to be the whole $\mathcal{J}\times\mathcal{J}$, but this would result in a significant number of redundant variables. At the description of the parameter $d_{\max}$ we will introduce a way how to initially reduce the number of train pairs potentially involved in any constraint; essentially this will be the set of trains that potentially “interact”. This set will be divided into subsets related to specific “interactions” in what follows.
  • The set $\mathcal{J}^{2\,(\text{single})}\subset\mathcal{J}^{2\,(\text{close})}$ contains all train pairs that can potentially meet and pass on a single-track line segment. This set can be obtained by considering all pairs of trains that are heading in the opposite direction on any single-track line, assuming all possible delays. If their train paths possibly intersect on any single-track segment, they are in $\mathcal{J}^{2\,(\text{single})}$.
  • The set $\mathcal{J}^{2\,(\text{headway})}\subset\mathcal{J}^{2\,(\text{close})}$ contains all train pairs that can violate the minimal headway condition, that is, those which potentially follow each other heading in the same direction, assuming any of the possible delays that can occur in the model instance.
  • The sets $\mathcal{J}^{2\,(\text{turn})}_{s}\subset\mathcal{J}\times\mathcal{J},\ \left(\forall s\in\mathcal{S}\right)$ are the sets of all pairs of trains so that the first train of the pair terminates at station $s$ and its rolling stock continues as the second train of the pair. These sets are deduced from the rolling stock circulation plan and the timetable.
  • The sets $\mathcal{J}^{2\,(\text{track})}_{s}\subset\mathcal{J}^{2\,(\text{close})},\ \left(\forall s\in\mathcal{S}\right)$ are the sets of train pairs that can possibly occupy the same track on the station $s$, thereby competing for the same track.
  • The sets $\mathcal{J}^{2\,(\text{switch},\text{out})}_{s}\subset\mathcal{J}^{2\,(\text{close})},\ \left(\forall s\in\mathcal{S}\right)$ are the sets of train pairs that can possibly compete for the same interlocking area of $s$ upon their departure from station $s$.
  • The sets $\mathcal{J}^{2\,(\text{switch},\text{out},\text{in})}_{s,s^{\prime}}\subset\mathcal{J}^{2\,(\text{close})},\ \left(\forall(s,s^{\prime})\in\mathcal{S}^{\times 2}\right)$ are the sets of train pairs that can possibly compete for the same interlocking area of $s$ in order to have $j$ to depart from $s$ while $j^{\prime}$ arrive at $s$ from the direction of $s^{\prime}$. For these train pairs, the respective constraints will ensure that they do not pass the same interlocking area at the same time upon departure.
  • The sets $\mathcal{J}^{2\,(\text{switch},\text{in},\text{noMO})}_{s,s^{\prime}}\subset\mathcal{J}^{2\,(\text{close})},\ \left(\forall(s,s^{\prime})\in\mathcal{S}^{\times 2}\right)$ are the sets of train pairs that can possibly compete for the same interlocking area of $s$ upon their arrival at $s$ from the direction of $s^{\prime}$, and there is no M-O possibility for them between $s$ and $s^{\prime}$.
  • The sets $\mathcal{J}^{2\,(\text{switch},\text{in},\text{MO})}_{s,s^{\prime}}\subset\mathcal{J}^{2\,(\text{close})},\ \left(\forall(s,s^{\prime})\in\mathcal{S}^{\times 2}\right)$ are the sets of train pairs that can possibly compete for the same interlocking area of $s$ upon their arrival at $s$ so that either both of them come from the direction of $s^{\prime}$ but there is a M-O possibility for them between $s$ and $s^{\prime}$, or one of them is approaching $s$ from a direction other than $s^{\prime}$. The constraints related to these sets and those related to $\mathcal{J}^{2\,(\text{switch},\text{in},\text{MO})}_{s,s^{\prime}}$ together will ensure that pairs of inbound trains do not use the same interlocking area at the same time upon arrival, which is also enforced by the interlocking system. The possibility for M-O between the current and previous decision stations (e.g. using multiple tracks) is to be modeled specifically, hence the distinction between these sets and the $\mathcal{J}^{2\,(\text{switch},\text{in},\text{noMO})}_{s,s^{\prime}}$-s.
  • The sets $\mathcal{C}^{2}_{j},\ \left(\forall j\in\mathcal{J}\right)$ are the sets of all station pairs that are subsequent in the route of $j$.
  • The set $\mathcal{C}^{2\,(\text{common})}_{j,j^{\prime}},\ \left(\forall(j,j^{\prime})\in\mathcal{J}^{\times 2}\right)$ are the sets of all station pairs that appear as subsequent stations in the route of both $j$ and $j^{\prime}$ heading in the same direction.
  • The sets $\mathcal{C}^{2\,(\text{common, single})}_{j,j^{\prime}},\ \left(\forall(j,j^{\prime})\in\mathcal{J}^{\times 2}\right)$ are those of all station pairs that appear as subsequent stations in the route of both $j$ and $j^{\prime}$ which are connected with a single-track line segment, and trains are heading in opposite directions. The order within the pairs is determined by $j$.
All these sets can be enumerated based on the input data in a straightforward manner. A brief summary of the sets is provided in Table 1.
4.1.2 Parameters
$\tau^{(\text{pass})}(j,s\rightarrow s^{\prime})$ running time of $j$ from $s$ to $s^{\prime}$
$\tau^{\text{(headway)}}(j,j^{\prime},s\rightarrow s^{\prime})$ minimal headway time for $j^{\prime}$ following $j$ from $s$ to $s^{\prime}$
$\tau^{(\text{dwell})}(s,j)$ minimal dwell time of $j$ at $s$
$\tau^{\text{(turn)}}(s,j,j^{\prime})$ rolling stock turnaround time for $j$ to continue as $j^{\prime}$ from s
$\sigma(j,s)$ originally scheduled departure of $j$ from $s$
$\upsilon(j,s)$ earliest possible factual departure time of $j$ from $s$
$d_{\text{max}}$ upper bound for secondary delays
Table 2: Summary of parameters
@@ 4.1.2 Parameters
The parameters that appear in our model, also summarized in Table 2, are the following:
@@
  • $\tau^{(\text{switch})}(j,j^{\prime},s)$ the time for which train $j$ occupies an interlocking area at $s$; this time should elapse before $j^{\prime}$ can start occupying the same area. Our model is simplified in this respect: we consider an upper estimate of this time valid for any switching zone of the station. The assumed values are determined from the parameters of station technology and the characteristics of the train pairs.
@@
  • $\sigma(j,s)$ is the scheduled departure time of $s$ from $j$, according to the original timetable.
  • $\upsilon(j,s)$ is the earliest possible departure time of $j$ from $s$ when a disturbance in the network happens. It is the maximum of the planned departure time $\sigma(j,s)$, and the technically feasible earliest departure time. The latter is calculated using the imaginary situation that the train starts from the initial location at the initial time defined by the initial delays (from the problem instance), and it moves to station $s$ as fast as technically possible, without having any other trains on the network, and then departs from there after the minimal stopping time has elapsed. These parameters are calculated before the optimization for the problem instances, based on the technological data of the network (minimum passing times) and the initial delays.
@@
  • $d_{\text{max}}$ is an upper bound assumed for the secondary delay $t^{(\text{out})}(j,s)-\upsilon(j,s)$. This parameter sets an upper limit for the possible secondary delays. Setting such a bound is not uncommon in the literature, see e.g. the work of [10]. We set the same value of this parameter for all our testing instances. It has to be big enough so that no secondary delay exceeds it (e.g. $d_{\max}=40$ excludes the possibility of a one-hour delay due to waiting for another delayed train). Meanwhile, it is desirable to set its value as small as possible; restricting the time variables to small intervals decreases the number of binary decision variables in the model.
@@
As mentioned in Section 4.1.1, the parameter $d_{\max}$ makes it possible to define $\mathcal{J}^{2\,(\text{close})}$, the set of trains potentially appearing in constraints, given a problem instance. Namely, a pair of trains $(j,j^{\prime})$ is included in this set if and only if they can meet at any station, given the original timetable, the initial delays, and assuming that no train can have a secondary delay greater than $d_{\max}$. The choice of a small $d_{\max}$ parameter results in a decrease of the model size, which is achieved via defining the set $\mathcal{J}^{2\,(\text{close})}$ with smaller cardinality. Choosing too small $d_{\max}$ can result in the infeasibility of the model.
4.1.3 Decision variables
$t^{(\text{out})}(j,s)$ departure time of train $j$ from decision station $s$
$y^{(\text{out})}(j,j^{\prime},s)$ order of trains $j$ and $j^{\prime}$ upon their departure from $s$
$y^{(\text{in})}(j,j^{\prime},s)$ order of trains $j$ and $j^{\prime}$ upon their arrival from $s$
$z(j,j^{\prime},s,s^{\prime})$ the order of trains at a resource such as e.g. a single-track segment, located between $s$ and $s^{\prime}$ the trains $j$ and $j^{\prime}$ are competing for.
Table 3: Summary of decision variables
@@ 4.1.3 Decision variables
Our decision variables, summarized in Table 3, are the following. First, we use departure time variables:
$t^{(\text{out})}(j,s)\in\mathbb{N},\ \ \ \text{where}\ \ \ \upsilon(j,s)\leq t^{(\text{out})}(j,s)\leq\upsilon(j,s)+d_{\max},$ (1)
@@
defining the departure time of train $j\in\mathcal{J}$ from station $s\in\mathcal{S}_{j}$. Such a variable is defined for all decision stations $\mathcal{S}_{j}$. The bounds follow trivially from the notion of the $\upsilon(j,s)$ and $d_{\text{max}}$ parameters. In the case of trains that terminate within the modeled part of the network, the last station is not included in $\mathcal{S}_{j}$. For the sake of notational convenience, we also introduce the arrival time of the trains at stations, $t^{(\text{in})}(j,s^{\prime})\in\mathbb{N}$. Having assumed fixed running times (green way policy), arrival times are defined by the previous departure times, (c.f. Equation (3)). Hence, arrival times are not independent decision variables themselves. This explains why we need sets and decision variables related to various combinations of inbound and outbound trains when describing conflicts in switching zones.
@@
In addition to the time variables, we use three sets of binary precedence variables. The decision variables $y^{(\text{out})}(j,j^{\prime},s)\in\{0,1\}$ determine the order of trains to leave stations: such a variable takes the value $1$ if train $j$ leaves station $s$ before train $j^{\prime}$, and it is 0 otherwise. The $(j,j^{\prime},s)$ tuples for which an $y$ variable is defined will be specified later. Similarly, the precedence variables $y^{(\text{in})}(j,j^{\prime},s)\in\{0,1\}$ prescribe the order at the entry to stations; the value is $1$ if $j$ arrives to $s$ before $j^{\prime}$.
The last set of precedence variables describes the precedence of trains $j$ and $j^{\prime}$ competing for a resource located between stations $s$ and $s^{\prime}$. The binary variable $z(j,j^{\prime},s,s^{\prime})\in\{0,1\}$ will be $1$ if $j$ uses the given resource before $j^{\prime}$. The quadruples $(j,j^{\prime},s,s^{\prime})$ for which we have such a variable will be specified later, but let us provide an example here. Consider two stations $s$ and $s^{\prime}$ that are connected by a single-track line segment, and trains $j$ and $j^{\prime}$ are heading in opposite directions, competing for the single-track segment as a resource. In this case, the variable describes which of the trains can occupy the single-track link first. (Recall that M-P is possible on decision stations only.)
@@ 4.2 ILP formulation
4.2.1 Objective function
Our goal is to minimize the secondary delays that occur in the analyzed part of the railway network. A suitable objective function is the weighted sum of secondary delays at the destination station:
$f(t)=\sum_{j\in\mathcal{J}}w_{j}\left(t^{(\text{out})}(j,s^{*})-\upsilon(j,s^{*})\right),$ (2)
where $s^{*}$ is the last element of $S_{j}$, and $w_{j}$ are weights for each train representing its priority. The weights have to be chosen in such a way that they reflect the rules and practices of the local dispatching system. As these rules are often formulated only qualitatively, the particular values are somewhat arbitrary; the model can be fine-tuned in this respect. (In Section 6.1 we provide more details of the weight assignments for Polish PKP PLK.) The reason for restricting ourselves to the secondary delays at the destination stations only is that this is the actual figure of merit chosen by the respective railway authorities; the generalization to other linear objectives such as, e.g., the weighted sum of secondary delays on all stations, is straightforward.
4.2.2 Constraints
@@
$t^{(\text{in})}(j,s^{\prime})=t^{(\text{out})}(j,s)+\tau^{(\text{pass})}(j,s\rightarrow s^{\prime}),\quad\forall j\in\mathcal{J},\ \forall(s,s^{\prime})\in\mathcal{C}_{j}.$ (3)
Recall that this condition represents the green way policy and makes the variables $t^{(\text{in})}$ depend trivially on the $t^{(\text{out})}$-s. The reason for keeping $t^{(\text{in})}$-s is that they make the model description more transparent.
@@
$t^{(\text{out})}(j^{\prime},s)\geq t^{(\text{out})}(j,s)+\tau^{(\text{headway})}(j,j^{\prime},s\rightarrow s^{\prime})-C\cdot y^{(\text{out})}(j^{\prime},j,s),$
$\forall(j,j^{\prime})\in\mathcal{J}^{2\,(\text{headway})},\ \forall(s,s^{\prime})\in\mathcal{C}^{2\,(\text{common})}_{j,j^{\prime}},$ (4)
where $C$ is a constant big enough to make the constraint satisfied whenever the binary variable $y^{(\text{out})}(j^{\prime},j,s)$ takes the value of $1$: in this case, $j^{\prime}$ goes first, so the constraint should be trivially satisfied. Otherwise speaking, the constraint should only be active for $y^{(\text{out})}(j^{\prime},j,s)=0$. (Constraints of this structure are often termed as big-M constraints in this context.) In order to find the smallest value of $C$ for this to hold, we can exploit the bounds on the time variables in Equation (1)): we have
$t^{(\text{out})}(j^{\prime},s)\geq\upsilon(j^{\prime},s)$ (5)
and
$t^{(\text{out})}(j,s)\leq\upsilon(j,s)+d_{\text{max}},$ (6)
hence, setting
$C=-\upsilon(j^{\prime},s)+\upsilon(j,s)+d_{\max}+\tau^{(\text{headway})}(j,j^{\prime},s\rightarrow s^{\prime}),$ (7)
if $y^{(\text{out})}(j^{\prime},j,s)=1$ then Equation (4) will take the form
$t^{(\text{out})}(j^{\prime},s)-\upsilon(j^{\prime},s)\geq t^{(\text{out})}(j,s)-d_{\text{max}}-\upsilon(j,s),$ (8)
which always holds as desired. Therefore, $C$ will be set according to Equation (7).
@@
$t^{(\text{out})}(j^{\prime},s^{\prime})\geq t^{\text{(in)}}(j,s^{\prime})-C\cdot z(j^{\prime},j,s^{\prime},s),$
$\forall(j,j^{\prime})\in\mathcal{J}^{2\,(\text{single})},\ \forall(s,s^{\prime})\in\mathcal{C}^{2\,(\text{common, single})}_{j,j^{\prime}},$ (9)
where $C$ is a big enough constant chosen similarly to that in Equation (4).
@@
$t^{(\text{out})}(j,s)\geq t^{(\text{in})}(j,s)+\tau^{(\text{dwell})}(j,s),\ \forall j\in\mathcal{J},\ \forall s\in S_{j}.$ (10)
@@
$t^{(\text{in})}(j^{\prime},s)\geq t^{(\text{out})}(j,s)-C\cdot y^{(\text{out})}(j^{\prime},j,s),\ {\forall s\in S,\ \forall(j,j^{\prime})\in\mathcal{J}^{2\,(\text{track})}_{s}},\$ (11)
where $C$ is chosen similarly to that in Equation (4) again. Note that this requirement may not be needed for depot tracks; the exceptions can be handled by the proper definition of $\mathcal{J}^{2\,(\text{track})}_{s}$.
@@
$t^{(\text{out})}(j^{\prime},s)\geq t^{(\text{out})}(j,s)+\tau^{(\text{switch})}(j,j^{\prime},s)-C\cdot y^{(\text{out})}(j^{\prime},j,s),$
$\forall s\in S,\ \forall(j,j^{\prime})\in\mathcal{J}^{2\,(\text{switch},\text{out})}_{s};$ (12)
$t^{(\text{in})}(j^{\prime},s)\geq t^{(\text{out})}(j,s)+\tau^{(\text{switch})}(j,j^{\prime},s)-C\cdot z(j^{\prime},j,s^{\prime},s),$
$\forall(s,s^{\prime})\in S^{\times 2},\ \forall(j,j^{\prime})\in\mathcal{J}^{2\,(\text{switch},\text{out},\text{in})}_{s,s^{\prime}};$ (13)
$t^{(\text{in})}(j^{\prime},s)\geq t^{(\text{in})}(j,s)+\tau^{(\text{switch})}(j,j^{\prime},s)-C\cdot y^{(\text{out})}(j^{\prime},j,s^{\prime}),$
$\forall(s,s^{\prime})\in S^{\times 2},\ \forall(j,j^{\prime})\in\mathcal{J}^{2\,(\text{switch},\text{in},\text{noMO})}_{s,s^{\prime}};$ (14)
$t^{(\text{in})}(j^{\prime},s)\geq t^{(\text{in})}(j,s)+\tau^{(\text{switch})}(j,j^{\prime},s)-C\cdot y^{(\text{in})}(j^{\prime},j,s),$
$\forall(s,s^{\prime})\in S^{\times 2},\ \forall(j,j^{\prime})\in\mathcal{J}^{2\,(\text{switch},\text{in},\text{MO})}_{s,s^{\prime}},$ (15)
with a choice of $C$ similar again to that in Equation (4). These constraints are also necessary for to make the realization of the modified timetable possible using the interlocking systems.